solve the following systems of equations.
x+4y=8(1)
x=3-4y(2)
What is the solution of the method?

Answer by Alan3354(69443)

solve the following systems of equations.
x+4y=8(1)
x=3-4y(2) --> x+4y=3
----------
x+4y=8
x+4y=3
-------- Subtract
0 = 5
--> inconsistent, no solution